Using the change in Westem G&Es EVA as the decision criterion, which type of investment recommendation should you make to your clients? Company Growth and Performance Metrics Percentage Change Metric Year 2 Year 1 General Metrics A hold recommendation A sell recommendation A buy recommendation Sales Net income Net cash flow (NCF) Net operating working capital (NOWC) Earnings per share (EPS) Dividends per share (DPS) Book value per share (BVPS) Cash flow per share (CFPS) Market price per share $8,400,000 $8,000,000 $576,000 $856,000 $680,400 Which of the following statements are correct? Check all that apply $3,719,250 $1.08 The percentage change in Western G&Es EVA $0.57 indicates that management has decreased its value Westem G&Es NCF is calculated by adding its annual interest expense to the corresponding years net income For any given year, one way to compute Western G&Es EVA is as the difference between its NOPAT (such as $696,000) and the product of its operating capital ($5,719,000) and its weighted average cost of capital ($7.30) $5.00 0.00% -15.53% $20.74 $19.75 MVA Calculation 41.77% | Market value of equity Book value of equity Market Value Added (MVA) $3,591,000 $2,660,000 $7,847,000 EVA Calculation Net operating profit after-tax (NOPAT) Investor-supplied operating capital Weighted average cost of capital Dollar cost of capital Return on invested capital (ROIC) Economic Value Added (EVA) $831,600 The percentage change in Western G&Es MVA indicates that its management has increased the firms value 35.00% 7.98% 7.30% Investor-supplied operating capital is recorded as 47.58% accounts payable, accruals, and short-term investments -11.50% $215,406
Answer #1

Change in Percentages

sales= 40,000/8,000,000= 0.5%

In net income= 104,400/576,000= 18.125

net cash flows for 2018= 9,080,400

1. Option A is the right answer, hold recommendation is the correct

2. option C is right answer